# Yet Another Representation of Binary Decision Trees: A Mathematical Demonstration

### v7

@@ Abstract
A decision tree looks like a simple computational graph without cycles, where only the leaf nodes specify the output values and the non-terminals specify their tests or split conditions. From the numerical perspective, we express decision trees in the language of computational graph. We explicitly parameterize the test phase, traversal phase and prediction phase of decision trees based on the bitvectors of non-terminal nodes. As shown, the decision tree is a shallow binary network in some sense. Especially, we introduce the bitvector matrix to implement the tree traversal in numerical approach, where the core is to convert the logical ‘AND’ operation to arithmetic operations. And we apply this numerical representation to extend and unify diverse decision trees in concept.
@@ 1.1 Motivation
In decision trees, non-terminal nodes apply a test function to the input sample, i.e., they ask a ‘yes-or-no’ question, such as the question ‘Is the first attribute is larger than 1?’. Then they lead the input sample to the next node based on the answer until reaching the terminal node. Different inputs have different depths and test functions in decision trees. The test function depends on the ‘splitting criteria’ to optimize during the growth of decision trees. In vanilla binary decision tree, the test function does not change the inputs and terminal/leaf node takes the mode of its instances ground truth empirical distribution as its prediction.
@@ 1.2 Related Work and Contribution
The related topic with our attention is the integration of decision trees and deep learning. There are some schemes to leverage neural networks and decision trees such as [16, 18, 34, 39, 35, 12, 29, 11]. Analogous to the fact that network and tree are two special types of graph, we will show that decision trees and artificial neural networks are special types of computational graph.
@@
We summarize the main conceptional contributions to deeper understanding of decision trees as follows
@@ 2.1 A Parameterization of Decision Trees
The first step is to find the false nodes in the decision tree with respect to the input $x\in\mathbb{R}^{n}$:
$h=\frac{\operatorname{Sgn}(Sx-t)+\vec{1}}{2}=D(Sx-t)$ (2)
where so-called ‘selection matrix’ $S\in\mathbb{R}^{n_{L}\times n}$ consists of one-hot row vector in $\mathbb{R}^{n}$ representing which feature is tested in a node; the elements of threshold vector $t\in\mathbb{R}^{n_{L}}$ are the optimal splitting points of each node associated with one feature; $\operatorname{Sgn}(\cdot)$ is the element-wise sign function and $\operatorname{Sgn}(0)=-1$;$D(x)$ is the binarized ReLU[8] function defined by
$\sigma(x_{i})=\begin{cases}1,&\text{if $x_{i}>0$}\\ 0,&\text{otherwise}\end{cases}.$
@@
$r=\operatorname{diag}(p)(Bh+\vec{1}),\quad i=\arg\max(r)$ (3)
where $\operatorname{diag}(\cdot)$ maps a vector to a diagonal matrix; $B\in\mathbb{B}^{L\times n_{L}}$ is the bitvector matrix of the decision tree; $p$ is a constant positive decreasing vector in $\mathbb{R}^{n_{L}}$ such as $(n_{L},n_{L}-1,\cdots,1)^{T}$ and $n_{L}$ is the number of the non-terminal nodes; $\arg\max(r)$ [^2] In general, the operator $\arg\max(r)$ is defined as $\arg\max(r)=\{j\mid r_{j}=\max(r)\}$ while here the maximum is unique.returns the minimum index of maximums in the vector $r$, i.e., $i=\arg\max(r)=\min\{j\mid r_{j}=\max(r)\}$. Every column of $B$ is a bit-vector of node; the matrix $Bh$ is the matrix multiplication product of matrix $B$ and $h$. Here $L$ is the number of terminal nodes(leaves). And the columns of true nodes will be multiplied by 0es and the columns of false nodes will keep unchanged. We call this step traversal phase.
@@
  • a binary decision tree with $S,t,v,B$
@@
3:procedure Scorer($\mathbf{x},v,B,S,t$)
4:     Compute the latent test vector $h=\frac{\operatorname{Sgn}(Sx-t)+\vec{1}}{2}$. $\triangleright$ test phase
5:     $r=\operatorname{diag}(p)(Bh+\vec{1})$.
6:     $j\leftarrow$ index of the maximum of $r$. $\triangleright$ traversal phase
@@
$\operatorname{diag}(p)[B\underbrace{\sigma(Sx-t)}_{\text{binary vector}}+\vec{1}]$
where if the test returns true, the element of the binary vector is zero; otherwise, the element of the binary vector is 1. The 1s select the bitvectors of false nodes and the 0s rule out the bitvectors of true nodes. If we change $1$ to any positive number, the first index of the maximums in the hidden state does not change. Thus binarized ReLU is not the unique option of activation function. For example, ReLU is an alternative to binarized ReLU in our representation. We will show more alternatives of binarized ReLU later. We can rewrite the decision tree in a more compact form:
$T(x)=v[\arg\max(\operatorname{diag}(p)[B\underbrace{\sigma(Sx-t)}_{\text{binary vector}}+\vec{1}])]$ (5)
where the notations are defined as above. And the consatnt vectors $p$ and $\vec{1}$ are used to ensure the maximum of the result is unique in (5), which are omitted based on our irregular $\arg\max$ operator definition in the rest of this paper.
Here is the hierarchical structure of decision tree in this case:
$\underbrace{x}_{\mathbb{R}^{n}}\xrightarrow{Sx-t}\underbrace{\tilde{h}}_{\mathbb{R}^{n_{L}}}\xrightarrow{\sigma(\tilde{h})}\underbrace{h}_{\mathbb{B}^{n_{L}}}\xrightarrow{Bh}\underbrace{b}_{\mathbb{Z}^{L}}\xrightarrow{i=\arg\max(b)}\underbrace{v[i]}_{\mathbb{O}}$
@@
  • the matrices $S$ and $B$ are binary, i.e., their elements are 0 or 1.
@@
All ‘if-then’ sentences are transformed to numerical computation. And the test phase and the traversal phase are split and separated. In test phase, we obtain the hidden state $h$ dependent on the selection matrix $S$ and threshold vector $t$: $h=\sigma(Sx-b)$. In traversal phase, we get the hidden state $H$ dependent on the bitvector matrix $B$: $l=Bh$ and then arrive at the terminal node by finding the first index of the maximums in the product: $v[i],i=\min\{j\mid l_{j}=\max(l)\}$. These phases play different roles while they are connected and associated. All the input samples share the same computational procedure. This makes it more computation intensive the in prediction. The good news is that we can apply the high performance technology such as fast matrix multiplication to decision tree.
The tuple $(S,t,B,v)$ may determine a decision tree and we propose the following statement:
When the processing order of the nodes are given, there is an one-to-one $M:T(n)\mapsto(S,t,B,v)$ where $T(n)$ is the set of all the decision trees where the number of non-terminal nodes is $n$; the tuple $(S,B,t,v)$ is defined as above.
@@
Based the this representation tuple $v,B,S,t$ of decision tree, we can identify every binary decision tree. All tree-based algorithms will benefit from such representation during prediction rather than training.
@@ 2.2 An Example of Binary Decision Trees
The threshold vector is $t=(1,4,3,2,5)^{T}$. The value vector is categorical or numerical, such as $v=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})^{T}$. The selection matrix $S$ and bit-vector matrix $B$ is
@@
[…]
@@
Suppose input vector $x=(2,1,2,2)^{T}$, its prediction is $v_{5}$ according to the figure 1. The computation procedure is as following
$\tilde{h}=Sx-t=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},h=\sigma(\tilde{h})=\begin{pmatrix}1\\ 0\\ 0\\ 0\\ 0\end{pmatrix},b=B\sigma(h)=\begin{pmatrix}0\\ 0\\ 0\\ 0\\ 1\\ 1\end{pmatrix},$
$i=\arg\max(b)=5,v[i]=v_{5},$
thus then the output is given by $T(x)=v_{5}$.
@@ 2.3 Categorical Features
$T(x)=v[\arg\max(B\,\sigma(Sx-t))]$
where $S$ is the one-hot row vector matrix and $B$ is the bitvector column matrix. The threshold value of categorical feature is also limited in the embedded (discrete) space. For example, if the categorical variables are stored in the format of string in computer and we use the hash function to transform them to integers as embedding, then the embedded integers are legal for the axis-aligned mapping $Sx$. Inspired by [28], we can the Lagrange polynomials.
@@
$T(x)=v[\arg\max(B\,\sigma(f(x)))]\$
where $f(x)=(f_{1}(x),f_{2}(x),\cdots,f_{n_{L}}(x))^{T}$ and $f_{i}(x)$ is the dummy-variable regression.
@@ 3.1.1 The Feature Transformation Mapping
It is the matrix $S$ which is the difference of vanilla decision trees and the oblique decision trees. It is one-hot vector in the column space of selection matrices $S$ associated with vanilla decision trees while it is real column vector in the selection matrices $S$ of oblique decision trees. Oblivious decision trees specify the selection matrices $S$ if the number of leaf node is given. Our framework can include more hybrid decision trees. For example, Erwan Scornet, Alex Davies and Zoubin Ghahramani apply kernel tricks to tree-based algorithms. We can replace the univariate or multivariate linear functions $Sx-t$ with nonlinear transformation functions $f=(f_{1},\cdots,f_{n_{L}})^{T}$:
$T(x)=v[\arg\max(B\,\sigma(f(x)))].$ (6)
@@
As other supervised learning, the choice of feature transformation function is an important part of feature engineering. The mapping $\sigma(f(x))$ reflects our tests on the raw input sample thus $f$ is to digitalize our prior experience and knowledge on the task to learn in some sense.
@@ 3.1.2 The Activation Function
There are many activation functions in deep neural network such as ReLU[8], maxout parametric ReLU, SELU, Swish, Mish. There are a partial list of activation functions in deep neural network at Simon’s blog and a more comprehensive list of activation functions at Wikipedia. They are motivated initially to solve the vanishing gradient problem and boost the stability as well as the generalization ability. In contrast to deep learning, there is no worry about the gradient vanishing or exploding problem during decision tree growth. However, we cannot apply activation functions in deep learning to decision trees without verification. It is not hard to verify that ReLU is an eligible alternative to the binarized ReLU or unit step function in (5) as well as its variants $\tau(x)=\alpha\max\{0,x\}$(so-called positively scaled ReLU) and $\tau(x)=\max\{0,\operatorname{sgn}(x)x^{2}\}$(so-called rectified quadratic unit).
@@
$v[\arg\max(B\sigma(Sx-t))]=v[\arg\max(B\tau(Sx-t))].$
@@
The $\tau(x):\mathbb{R}\mapsto\mathbb{R}^{+}$ is positivity-augmented if it is positive when the input $x$ is positive; otherwise, it is 0.
It is used to augment the positive evidence or signal of the tests and rule out the false or negative samples. It maps the non-positive part of real number $R$ to the original point so it is a truncation type function. For example, the cumulative distribution function $\sigma$ is an positivity-augmented function if its support is $(0,\infty)$ so that if $x\leq 0$ then $\sigma(x)=0$; otherwise $\sigma(x)>0$ and $\lim_{x\to\infty}\sigma(x)=1$. It is easy to adapt the penalty function such as MC+ for variable selection into positivity-augmented function. The penalty function $p$ is usually symmetric and non-negative, i.e.,
  • $p(x)=p(-x)$ and $p(0)=0$;
  • $p(x)\geq 0$.
If $p$ holds both properties, it is simply to adapt $p$ to a positivity-augmented function via $p^{+}=\max\{0,\operatorname{sgn}(x)p(x)\}$ where $\operatorname{sgn}$ is the sign function. It is also possible to modify odd functions to positivity-augmented functions such as $\tau(x)=\max\{0,x+\sin(x)\}$ and $\tau(x)=max\{0,x^{3}\}$. In fact, we can construct more general positivity-augmented functions such as $\tau(x)=\max\{0,\min\{x,\frac{1}{x}\}\}$, $\tau(x)=\max\{0,\frac{1}{|\log(x)|}\}$.
@@
$T(x)=v[\arg\max(Bh)],\\ h_{i}=\tau_{i}([(Sx-t)]_{i})$
where $[(Sx-t)]_{i}$ is the $i$th element of the vector $(Sx-t)$.

### v8

@@ Abstract
A decision tree looks like a simple directed acyclic computational graph, where only the leaf nodes specify the output values and the non-terminals specify their tests or split conditions. From the numerical perspective, we express decision trees in the language of computational graph. We explicitly parameterize the test phase, traversal phase and prediction phase of decision trees based on the bitvectors of non-terminal nodes. As shown, the decision tree is a shallow binary network in some sense. Especially, we introduce the bitvector matrix to implement the tree traversal in numerical approach, where the core is to convert the logical ‘AND’ operation to arithmetic operations. And we apply this numerical representation to extend and unify diverse decision trees in concept.
@@ 1.1 Motivation
In decision trees, non-terminal nodes apply a Boolean test function to the input sample, i.e., they ask a ‘yes-or-no’ question, such as the question ‘Is the first attribute is larger than 1?’. Then they lead the input sample to the next node based on the answer until reaching the terminal node. Different inputs have different depths and test functions in decision trees. The test function depends on the ‘splitting criteria’ to optimize during the growth of decision trees. In vanilla binary decision tree, the test function does not change the inputs and terminal/leaf node takes the mode of its instances ground truth empirical distribution as its prediction.
@@ 1.2 Related Work and Contribution
The related topic with our attention is the integration of decision trees and deep learning. There are some schemes to leverage neural networks and decision trees such as [15, 17, 30, 34, 31, 11, 26, 10].
Analogous to the fact that network and tree are two special types of graph, we will show that decision trees and artificial neural networks are special types of computational graph.
@@
Our main conceptional contributions includes:
@@ 2.1 A Parameterization of Decision Trees
The first step is to find the false nodes in the decision tree with respect to the input $\mathbf{x}\in\mathbb{R}^{n}$:
$\mathbf{h}=\frac{\operatorname{Sgn}(\mathrm{S}\mathbf{x}-\mathbf{t})+\vec{1}}{2}=\sigma(\mathrm{S}\mathbf{x}-\mathbf{t})$ (2)
where so-called ‘selection matrix’ $S\in\mathbb{R}^{n_{L}\times n}$ consists of one-hot row vector in $\mathbb{R}^{n}$ representing which feature is tested in a node; the elements of threshold vector $\mathbf{t}\in\mathbb{R}^{n_{L}}$ are the optimal splitting points of each node associated with one feature; $\operatorname{Sgn}(\cdot)$ is the element-wise sign function and $\operatorname{Sgn}(0)=-1$; $\sigma(x)$ is the binarized ReLU[7] function defined by
$\sigma(z)=\begin{cases}1,&\text{if $z>0$}\\ 0,&\text{otherwise}\end{cases}.$
@@
$\mathbf{p}=\mathrm{B}\mathbf{b}+\vec{1},\quad i=\arg\max(\mathbf{p})$ (3)
where the matrix $\mathrm{B}\in\mathbb{B}^{L\times n_{L}}$ is the bitvector matrix of the decision tree; the integer $n_{L}$ is the number of the non-terminal nodes; $\arg\max(\mathbf{p})$ [^2] Note that the operator $\arg\max(\mathbf{p})$ generally returns all indices of the max element of the array $p$, i.e., $\arg\max(\mathbf{p})=\{j\mid p_{j}=\max(\mathbf{p})\}$.returns the smallest index of maximums in the vector $p$, i.e., $i=\arg\max(\mathbf{p})=\min\{j\mid p_{j}=\max(\mathbf{p})\}$. Every column of $\mathrm{B}$ is a bit-vector of node; the matrix $\mathrm{B}\mathbf{b}$ is the matrix multiplication product of matrix $\mathrm{B}$ and $\mathbf{h}$. Here $L$ is the number of terminal nodes(leaves). And the columns of true nodes will be multiplied by 0es and the columns of false nodes will keep unchanged. We call this step traversal phase.
@@
  • a binary decision tree with $v,\mathrm{B},\mathrm{S},\mathbf{t}$
@@
3:procedure Scorer($\mathbf{x},v,\mathrm{B},\mathrm{S},\mathbf{t}$)
4:     Compute the latent test vector $\mathbf{h}=\frac{\operatorname{Sgn}(\mathrm{S}\mathbf{x}-\mathbf{t})+\vec{1}}{2}$. $\triangleright$ test phase
5:     $\mathbf{p}=\mathrm{B}\mathbf{b}+\vec{1}$.
6:     $j\leftarrow$ the smallest index of the maximum of $\mathbf{p}$. $\triangleright$ Traversal phase
@@
$[\mathrm{B}\underbrace{\sigma(\mathrm{S}\mathbf{x}-\mathbf{y})}_{\text{binary vector}}+\vec{1}]$
where if the test returns true, the element of the binary vector is zero; otherwise, the element of the binary vector is 1. The 1s select the bitvectors of false nodes and the 0s rule out the bitvectors of true nodes. If we change $1$ to any positive number, the first index of the maximums in the hidden state does not change. Thus, binarized ReLU is not the unique option of activation function. For example, ReLU is an alternative to binarized ReLU in our representation. We will show more alternatives of binarized ReLU later. We can rewrite the decision tree in a more compact form:
$T(\mathbf{x})=v[\arg\max([\mathrm{B}\underbrace{\sigma(\mathrm{S}\mathbf{x}-\mathbf{y})}_{\text{binary vector}}+\vec{1}])]$ (5)
where the notations are defined as above. And the consatnt vectors $p$ and $\vec{1}$ are used to ensure the maximum of the result is unique in (5), which are omitted based on our irregular $\arg\max(\cdot)$ operator definition in the rest of this paper.
Here is the data flow of decision tree in this case:
$\underbrace{\mathbf{x}}_{\mathbb{R}^{n}}\xrightarrow{\mathrm{S}\mathbf{x}-\mathbf{y}}\underbrace{\tilde{h}}_{\mathbb{R}^{n_{L}}}\xrightarrow{\sigma(\tilde{\mathbf{h}})}\underbrace{\mathbf{h}}_{\mathbb{B}^{n_{L}}}\xrightarrow{\mathrm{B}\mathbf{b}}\underbrace{\mathbf{p}}_{\mathbb{Z}^{L}}\xrightarrow{i=\arg\max(\mathbf{b})}\underbrace{v[i]}_{\mathbb{O}}$
@@
  • the matrices $\mathrm{S}$ and $\mathrm{B}$ are binary, i.e., their elements are 0 or 1.
@@
All ‘if-then’ sentences are transformed to numerical computation. And the test phase and the traversal phase are split and separated. In test phase, we obtain the hidden state $\mathbf{h}$ dependent on the selection matrix $\mathrm{S}$ and threshold vector $\mathbf{t}$: $\mathbf{h}=\sigma(\mathrm{S}\mathbf{x}-\mathbf{b})$. In traversal phase, we get the hidden state $\mathbf{p}$ dependent on the bitvector matrix $\mathrm{B}$: $\mathbf{p}=\mathrm{B}\mathbf{b}$ and then arrive at the terminal node by finding the first index of the maximums in the product: $v[i],i=\min\{j\mid l_{j}=\max(l)\}$. These phases play different roles while they are connected and associated. All the input samples share the same computational procedure. This makes it more computation intensive in prediction. The good news is that we can apply the high performance techniques such as fast matrix multiplication to decision trees.
The tuple $(v,\mathrm{B},\mathrm{S},\mathbf{t})$ may determine a decision tree and we propose the following statement:
When the processing order of the nodes are given, there is an one-to-one $M:T(n)\mapsto(v,\mathrm{B},\mathrm{S},\mathbf{t})$ where $T(n)$ is the set of all the decision trees where the number of non-terminal nodes is $n$; the tuple $(v,\mathrm{B},\mathrm{S},\mathbf{t})$ is defined as above.
@@
Based the this representation tuple $v,\mathrm{B},\mathrm{S},\mathbf{t}$ of decision tree, we can identify every binary decision tree. All tree-based algorithms will benefit from such representation during prediction rather than training.
@@ 2.2 An Example of Binary Decision Trees
The threshold vector is $\mathbf{t}=(1,4,3,2,5)^{T}$. The value vector is categorical or numerical, such as $\mathbf{v}=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})^{T}$. The selection matrix $\mathrm{S}$ and bit-vector matrix $\mathrm{B}$ is
@@
$\mathbf{x}[1]\leq 1?$ $\mathbf{x}[2]\leq 4?$$\mathbf{x}[3]\leq 3?$$\mathbf{x}[2]\leq 2?$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$\mathbf{x}[4]\leq 5?$$v_{2}$$v_{3}$rueFalseTrueFalseTrueFalseTrueFalseTrueFalse
@@
Suppose input vector $\mathbf{x}=(2,1,2,2)^{T}$, its prediction is $v_{5}$ according to the figure 1. The computation procedure is as following
$\tilde{h}=\mathrm{S}\mathbf{x}-\mathbf{t}=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},\mathbf{h}=\sigma(\tilde{h})=\begin{pmatrix}1\\ 0\\ 0\\ 0\\ 0\end{pmatrix},\mathbf{p}=\mathrm{B}\sigma(\mathbf{h})=\begin{pmatrix}0\\ 0\\ 0\\ 0\\ 1\\ 1\end{pmatrix},$
$i=\arg\max(\mathbf{p})=5,v[i]=v_{5},$
thus then the output is given by $T(\mathbf{x})=v_{5}$.
@@ 2.3 Categorical Features
$T(\mathbf{x})=v[\arg\max(\mathrm{B}\,\sigma(\mathrm{S}\mathbf{x}-\mathbf{y}))]$
where $\mathrm{S}$ is the one-hot row vector matrix and $\mathrm{B}$ is the bitvector column matrix. The threshold value of categorical feature is also limited in the embedded (discrete) space. For example, if the categorical variables are stored in the format of string in computer and we use the hash function to transform them to integers as embedding, then the embedded integers are legal for the axis-aligned mapping $\mathrm{S}\mathbf{x}$. Inspired by [25], we can the Lagrange polynomials.
@@
$T(\mathbf{x})=v[\arg\max(\mathrm{B}\,\sigma(f(\mathbf{x})))]\$
where $f(\mathbf{x})=(f_{1}(\mathbf{x}),f_{2}(\mathbf{x}),\cdots,f_{n_{L}}(\mathbf{x}))^{T}$ and $f_{i}(\mathbf{x})$ is the dummy-variable regression.
@@ 3.1.1 The Feature Transformation Mapping
It is the matrix $\mathrm{S}$ which is the difference of vanilla decision trees and the oblique decision trees. It is one-hot vector in the column space of selection matrices $\mathrm{S}$ associated with vanilla decision trees while it is real column vector in the selection matrices $\mathrm{S}$ of oblique decision trees. Oblivious decision trees specify the selection matrices $\mathrm{S}$ if the number of leaf node is given. Our framework can include more hybrid decision trees. For example, Erwan Scornet, Alex Davies and Zoubin Ghahramani apply kernel tricks to tree-based algorithms. We can replace the univariate or multivariate linear functions $\mathrm{S}\mathbf{x}-\mathbf{t}$ with nonlinear transformation functions $f=(f_{1},\cdots,f_{n_{L}})^{T}$:
$T(\mathbf{x})=v[\arg\max(\mathrm{B}\,\sigma(f(\mathbf{x})))].$ (6)
@@
As other supervised learning, the choice of feature transformation function is an important part of feature engineering. The mapping $\sigma(f(\mathbf{x}))$ reflects our tests on the raw input sample thus $f$ is to digitalize our prior experience and knowledge on the task to learn in some sense.
@@ 3.1.2 The Activation Function
There are many activation functions in deep neural network such as ReLU[7], maxout parametric ReLU, SELU, Swish, Mish. There are a partial list of activation functions in deep neural network at Simon’s blog and a more comprehensive list of activation functions at Wikipedia. They are motivated initially to solve the vanishing gradient problem and boost the stability as well as the generalization ability. In contrast to deep learning, there is no worry about the gradient vanishing or exploding problem during decision tree growth. However, we cannot apply activation functions in deep learning to decision trees without verification. It is not hard to verify that ReLU is an eligible alternative to the binarized ReLU or unit step function in (5) as well as its variants $\tau(z)=\alpha\max\{0,z\}$(so-called positively scaled ReLU) and $\tau(z)=\max\{0,\operatorname{sgn}(z)z^{2}\}$(so-called rectified quadratic unit).
@@
$v[\arg\max(\mathrm{B}\sigma(\mathrm{S}\mathbf{x}-\mathbf{t}))]=v[\arg\max(\mathrm{B}\tau(\mathrm{S}\mathbf{x}-\mathbf{t}))].$
@@
The $\tau(z):\mathbb{R}\mapsto\mathbb{R}^{+}$ is positivity-augmented if it is positive when the input $z$ is positive otherwise 0.
It is used to augment the positive evidence or signal of the tests and rule out the false or negative samples. It maps the non-positive part of real number $R$ to the original point so it is a truncation type function. For example, the cumulative distribution function $\sigma$ is an positivity-augmented function if its support is $(0,\infty)$ so that if $z\leq 0$ then $\sigma(z)=0$; otherwise $\sigma(z)>0$ and $\lim_{z\to\infty}\sigma(z)=1$. It is easy to adapt the penalty function such as MC+ for variable selection into positivity-augmented function. The penalty function $p$ is usually symmetric and non-negative, i.e.,
  • $p(z)=p(-z)$ and $p(0)=0$;
  • $p(z)\geq 0$.
If $p(\cdot)$ holds both properties, it is simply to adapt $p$ to a positivity-augmented function via $p^{+}=\max\{0,\operatorname{sgn}(x)p(x)\}$ where $\operatorname{sgn}$ is the sign function. It is also possible to modify odd functions to positivity-augmented functions such as $\tau(z)=\max\{0,z+\sin(z)\}$ and $\tau(z)=max\{0,z^{3}\}$. In fact, we can construct more general positivity-augmented functions such as $\tau(z)=\max\{0,\min\{z,\frac{1}{z}\}\}$, $\tau(z)=\max\{0,\frac{1}{|\log(z)|}\}$.
@@
$T(\mathbf{x})=v[\arg\max(\mathrm{B}\mathbf{b})],\\ h_{i}=\tau_{i}([(\mathrm{S}\mathbf{x}-\mathbf{t})]_{i})$
where $[(\mathrm{S}\mathbf{x}-\mathbf{t})]_{i}$ is the $i$th element of the vector $(\mathrm{S}\mathbf{x}-\mathbf{t})$.